pure acid is to be added to a 10% acid solution to obtain 90L of 42% solution what amounts of each should be used?

Answer by addingup(3677)

1x+0.10(90-x) = 0.42(90)
x+9-0.10x = 37.8
0.9x = 28.8
x = 32L of pure acid to 90-32 = 58L of 10% solution to get 90L at 42%